# Parameter-Efficient Fine-Tuning via Circular Convolution

### v1

@@ 1 Introduction
In recent years, Large Foundation Models (LFMs) have surged in popularity due to their remarkable performance across various tasks in natural language processing Brown et al. (2020); Touvron et al. (2023), computer vision Radford et al. (2021); Kirillov et al. (2023), and other domains Li et al. (2024). Characterized by their vast number of parameters and substantial computational demands, these models have set new benchmarks in accuracy and versatility. However, their size and complexity pose significant challenges for efficient fine-tuning, particularly in resource-constrained environments Malladi et al. (2023); Zhang et al. (2024). To address this issue, Parameter-efficient fine-tuning (PEFT) techniques Mangrulkar et al. (2022), such as Low-Rank Adaptation (LoRA) Hu et al. (2021), have emerged as practical solutions.
@@
where $\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d\times d}$ are weight matrices, $\mathbf{B}\in\mathbb{R}^{d\times r},\mathbf{A}\in\mathbb{R}^{r\times d}$ are low-rank matrices to construct $\Delta\mathbf{W}$ and $\mathbf{x}\in\mathbb{R}^{d}$ is the activations. The number of trainable parameters is $2rd$, prompting the choice $r\ll d$ (e.g., $r=8$ for $d=1024$) for achieving high parameter efficiency. However, as discussed in Zeng & Lee (2023), LoRA’s capacity to represent a target model is strictly limited by $r$. To address this trade-off between performance and efficiency, Kopiczko et al. (2023) proposed Vector Random Matrix Adaptation (VeRA), which achieves competitive performance with significantly fewer trainable parameters through fixed random-matrix projections. Despite its low parameter count, VeRA requires substantial computational power and memory due to the large size of the random matrices used in projection. This raises the following open research question in the context of PEFT:
@@
How can high-rank adaptation be achieved without incurring significant costs in terms of computational power, memory usage, and trainable parameters?
@@
As discussed in Dosovitskiy et al. (2020), dense linear layers lack strong inductive biases, limiting transformers’ ability to compete with Convolutional Neural Networks (CNNs) when data is scarce. Given the constraints of a small training set for the downstream task, we posit that incorporating a stronger inductive bias could enhance adaptation. In light of this, we propose Circular Convolution Adaptation (C3A). Circular convolution is famous in signal processing Li et al. (2020) and cryptography Dworkin et al. (2001). This operator can be equivalently represented as multiplication by a circulant matrix, offering flexibility in rank unconstrained by the number of trainable parameters. Moreover, leveraging the Fast Fourier Transform (FFT), C3A ensures better computational and memory efficiency than directly multiplying a delta matrix Bamieh (2018), rendering it competitive with LoRA from the efficiency aspect. We summarize our contribution as below:
@@
① We introduce C3A, a novel approach for PEFT using convolution. This method leverages circular convolution kernels and their equivalent circulant matrices to provide a flexible rank, which is free of linear constraint by the number of trainable parameters, for the delta matrix.
@@
④ We validate C3A through comprehensive fine-tuning experiments across diverse tasks including natural language understanding, instruction tuning, image classification, and text-to-image generation. Our experiments demonstrate C3A’s outstanding accuracy and memory merits compared to existing state-of-the-art methods.
@@ 2.1 Parameter-Efficient Fine-Tuning
Research on Parameter-Efficient Fine-Tuning (PEFT) has generally progressed along three main directions. The first direction involves partially updating the pre-trained neural network (e.g., the body Oh et al. (2020) or the biases Zaken et al. (2021)). Initial methods relied on hand-crafted heuristics Han et al. (2015); Oh et al. (2020) to identify which parameters are crucial and should be fine-tuned. More advanced approaches employ optimization techniques Guo et al. (2020); Xu et al. (2021); Fu et al. (2023). For example, Guo et al. Guo et al. (2020) reformulated the discrete optimization problem into a continuous one using Bernoulli masks and the Gumbel-softmax approximation.
The second direction emerged to maintain the integrity of the pre-trained model and enable a high degree of parameter sharing through adapter-based methods He et al. (2021); Rebuffi et al. (2017); Rücklé et al. (2020); Liu et al. (2022). These works focus on integrating additional modules, termed adapters, to fit the downstream task, effectively decoupling the pre-trained model parameters from those specific to the downstream task. Prompt Tuning Brown et al. (2020); Gao et al. (2020); Chen et al. (2023) and Prefix Tuning Li & Liang (2021) also fall into this category, despite ignoring potential semantic meanings.
The final direction is characterized by delta-weight-based methods, such as Low-Rank Adaptation (LoRA) Hu et al. (2021) and Orthogonal Fine-tuning (OFT) Qiu et al. (2023). These methods bridge the gap between the pre-trained model and the downstream task by adaptive delta weights, which is learned and stored separately while used in combination with the pre-trained weights. This unique design enables disentanglement and parameter sharing while preserving the ability of integrating without additional inference cost. LoRA models the delta-weights by an additive matrix while OFT does it by a multiplicative one. To further improve either parameter efficiency or performance, many variants has been proposed for both of the methods Kopiczko et al. (2023); Liu et al. (2024; 2023); Yuan et al. (2024). However, these methods can hardly achieve high parameter efficiency and performance without incurring heavy computation and memory usage.
@@ 2.2 Circular Convolution
Despite its success in small neural networks such as LeNet Cheng et al. (2015), circular convolution has not demonstrated lossless performance in modern large foundational models (LFMs) or even in their base architecture, the transformer. This limitation may be attributed to the conflict between its high intrinsic bias (i.e., the circulant pattern) and the vast amount of data required for training LFMs. Conversely, when fine-tuning LFMs, it is often impractical to collect as much data as needed for training from scratch. In such scenarios, the intrinsic bias of circular convolution could potentially serve as a regularization mechanism, thereby benefiting the fine-tuning process.
@@ 3 Method
Figure 1: Overview of LoRA (A) and our C3A (B,C) method. In LoRA, only low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ are trained and the delta weight is represented by their product (i.e., $\Delta\mathbf{W}=\mathbf{BA}$). The total trainable parameter number is $2rd$, which is assosiated with the rank of the delta weight. In C3A, circular convolution kernels $\mathbf{w}$ are tuned to adapt to the downstream task and the delta weight is represented by the (block-)circular matrix they construct (i.e., $\Delta\mathbf{W}=\mathcal{C}_{b}(\Delta\mathbf{w})$). The total trainable parameter number is $bd$, which disentangles with the rank of the delta weight.
@@ 3.1 Notations
We denote circular convolution operator as $\star$. The adapted weight matrix, the original weight matrix, and the delta matrix are denoted by $\mathbf{W}$, $\mathbf{W}_{0}$, and $\Delta\mathbf{W}$, respectively. The activation vector of the previous layer is denoted by $\mathbf{x}$. The post-transformation vector is $\mathbf{z}$, where $\mathbf{z}=\mathbf{W}\mathbf{x}$, and the incremental part is denoted by $\Delta\mathbf{z}$, where $\Delta\mathbf{z}=\Delta\mathbf{W}\mathbf{x}$. The length of the activation vector is $d$. The matrices $\mathbf{A}$ and $\mathbf{B}$ are low-rank matrices introduced by LoRA to represent $\Delta\mathbf{W}$, with $r$ being their rank. $r_{v}$ is used to specify the random projection matrix rank used in VeRA. The circular convolution kernel of C3A is denoted by $\Delta\mathbf{w}$. The loss function is represented by $\mathcal{L}$. The Fast Fourier Transform and its inverse are denoted by $\mathrm{FFT}$ and $\mathrm{iFFT}$, respectively. The Hadamard product is denoted by $\circ$. In terms of dimensions, we have $\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d\times d}$, and $\mathbf{x},\mathbf{z},\Delta\mathbf{z},\Delta\mathbf{w}\in\mathbb{R}^{d}$. The matrices $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d}$ are controlled by the rank $r$.
@@ 3.2 Circular Convolution
The circular convolution operator is defined as $\Delta\mathbf{z}=\Delta\mathbf{w}\star\mathbf{x}=\mathcal{C}(\Delta\mathbf{w})\mathbf{x}$, where $\mathcal{C}(\Delta\mathbf{w})$ is the circulant matrix of $\Delta\mathbf{w}$. The first row of $\mathcal{C}(\Delta\mathbf{w})$ is $\Delta\mathbf{w}$ and each row equals the row above it periodically shifted to right by one element. Concretely,
$\mathcal{C}(\Delta\mathbf{w})=\begin{bmatrix}\Delta w_{1}&\Delta w_{2}&\cdots&\Delta w_{d-1}&\Delta w_{d}\\ \Delta w_{d}&\Delta w_{1}&\cdots&\Delta w_{d-2}&\Delta w_{d-1}\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ \Delta w_{3}&\Delta w_{4}&\cdots&\Delta w_{1}&\Delta w_{2}\\ \Delta w_{2}&\Delta w_{3}&\cdots&\Delta w_{d}&\Delta w_{1}\\ \end{bmatrix}\in\mathbb{R}^{d\times d}.$
Theoretically, the rank of $\mathcal{C}(\Delta\mathbf{w})$ is $d-\mathrm{Deg}(\mathrm{gcd}(f(x),x^{d}-1))$ Ingleton (1956), where $\mathrm{Deg}(\cdot)$ represents the degree of a polynomial, $f(x)$ is the associated polynomial of $\Delta\mathbf{w}$ (i.e., $f(x)=\sum_{i=1}^{d}\Delta w_{i}x^{i-1}$), and $\mathrm{gcd}(\cdot)$ means the greatest common divisor. By learning $\Delta\mathbf{w}$ in the $\mathbb{R}^{n}$ oracle, C3A automatically achieves dynamic rank selection that is not linearly bounded by the number of learnable parameters as LoRA.
To achieve efficiency, we leverage the beautiful circulant structure of $\mathcal{C}(\Delta\mathbf{w})$, which makes it diagonalizable by the Fourier basis ($\mathbf{F}$) . In math, it can be described as $\mathcal{C}(\Delta\mathbf{w})=\mathbf{F}\frac{\Lambda}{d}\mathbf{F}^{-1}$ Golub & Van Loan (1996), where $\Lambda$ is its eigenvalues and can be calculated by a Fourier transform of the first row (i.e., $\Lambda=\mathrm{diag}(\mathbf{F}\Delta\mathbf{w})$). Therefore, we can calculate $\Delta\mathbf{w}\star\mathbf{x}$ as
@@ 3.3 Backpropagation
Since $\Delta\mathbf{z}=\mathcal{C}(\Delta\mathbf{w})\mathbf{x}$, it follows that $\frac{\partial\Delta\mathbf{z}}{\partial\mathbf{x}}=\mathcal{C}(\Delta\mathbf{w})$. For $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}$, we note that the circular convolution operation is commutative (i.e., $\Delta\mathbf{w}\star\mathbf{x}=\mathbf{x}\star\Delta\mathbf{w}$), thus $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}=\mathcal{C}(\mathbf{x})$. Substituting these results into Equation 2, we obtain:
@@ 3.4 Block-Circular Convolution
Despite the elegance and efficiency of the circular convolution operator, it is worth noting that we can not alter the number of learnable parameters due to the requirement of the convolution kernel being the same size as the activation vector. To enable this flexibility in case of difficult downstream tasks, we resort to the block-circular convolution. We can simply split the activation vector $\mathbf{x}$ (and the post-transformation vector $\mathbf{z}$) into blocks, and assign each pair of blocks with different convolution kernels. To be concrete,
$\mathbf{x}=\begin{bmatrix}\mathbf{x}_{1}&\mathbf{x}_{2}&\cdots&\mathbf{x}_{b}\end{bmatrix}\qquad\Delta\mathbf{z}=\begin{bmatrix}\Delta\mathbf{z}_{1}&\Delta\mathbf{z}_{2}&\cdots&\Delta\mathbf{z}_{b}\end{bmatrix},$
where $b$ is the block number. We will need $b^{2}$ convolution kernels to densely connect these blocks, which can be expressed as
$\Delta\mathbf{z}_{i}=\sum_{j=1}^{b}\Delta\mathbf{w}_{ij}\star\mathbf{x}_{j},i\in\{1,2,\cdots,b\},$
and can also be represented by a block-circular matrix
$\Delta\mathbf{z}=\mathcal{C}_{b}(\Delta\mathbf{w})\mathbf{x}\qquad\mathcal{C}_{b}(\Delta\mathbf{w})=\begin{bmatrix}\mathcal{C}(\Delta\mathbf{w}_{11})&\mathcal{C}(\Delta\mathbf{w}_{12})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{1b})\\ \mathcal{C}(\Delta\mathbf{w}_{21})&\mathcal{C}(\Delta\mathbf{w}_{22})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{2b})\\ \cdots&\cdots&\cdots&\cdots\\ \mathcal{C}(\Delta\mathbf{w}_{b1})&\mathcal{C}(\Delta\mathbf{w}_{b2})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{bb})\end{bmatrix}.$
In this case, $\Delta\mathbf{w}_{ij}\in\mathbb{R}^{\frac{d}{b}}$ (it is required that $d$ is divisible by $b$) and we have $b^{2}\frac{d}{b}=bd$ learnable parameters. $b$ acts as a hyperparameter controlling the number of learnable parameters here, just like $r$ for LoRA. However, we highlight that while $r$ controls the rank of the delta matrix and learnable parameter number at the same time, $b$ only affects the latter. Disentangling these attributes enables higher adaptation flexibility and potential better results.

### v2

@@ 1 Introduction
Figure 1: Relative comparison of C3A and baselines on RoBERTa-Base. The Pearson Correlation Coefficient (PCC) is evaluated on STS-B and the Matthew’s Correlation Coefficient (MCC) on CoLA. Accuracies across SST-2, MRPC, QNLI, and RTE are averaged and reported as Acc. #Param denotes the count of learnable parameters (excluding the head), while Memory indicates the peak GPU memory usage during training. Metrics highlighted in blue represent values where higher is better, and those in red indicate values where lower is better. Refer to Table 2 for absolute statistics.
In recent years, Large Foundation Models (LFMs) have surged in popularity due to their remarkable performance across various tasks in natural language processing (NLP) Brown et al. (2020); Touvron et al. (2023), computer vision (CV) Radford et al. (2021); Kirillov et al. (2023), and other domains Li et al. (2024). Characterized by their vast number of parameters and substantial computational demands, these models have set new benchmarks in accuracy and versatility. However, their size and complexity pose significant challenges for efficient fine-tuning, particularly in resource-constrained environments Malladi et al. (2023); Zhang et al. (2024b). To address this issue, Parameter-efficient fine-tuning (PEFT) techniques Mangrulkar et al. (2022), such as Low-Rank Adaptation (LoRA) Hu et al. (2021), have emerged as outstanding practical solutions.
@@
where $\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$ are weight matrices, $\mathbf{B}\in\mathbb{R}^{d_{1}\times r},\mathbf{A}\in\mathbb{R}^{r\times d_{2}}$ are low-rank matrices to construct $\Delta\mathbf{W}$ and $\mathbf{x}\in\mathbb{R}^{d_{2}}$ is the activations. The number of trainable parameters is $r(d_{1}+d_{2})$, prompting the choice $r\ll\min(d_{1},d_{2})$ (e.g., $r=8$ for $d_{1}=d_{2}=1024$) for achieving high parameter efficiency. However, as discussed in Zeng & Lee (2023), LoRA’s capacity to represent a target model is strictly limited by $r$. To address this trade-off between performance and efficiency, Kopiczko et al. (2023) proposed Vector Random Matrix Adaptation (VeRA), which achieves competitive performance with significantly fewer trainable parameters through fixed random-matrix projections. Despite its low parameter count, VeRA requires substantial computational power and memory due to the large size of the random matrices used for projection. This raises the following open research question in the context of PEFT:
@@
How can high-rank adaptation be achieved without incurring significant costs in terms of time complexity, memory usage, and trainable parameters?
@@
As discussed in Dosovitskiy et al. (2020), dense linear layers lack inductive biases, which poses challenge on the optimization process and therefore limits transformers’ ability to compete with Convolutional Neural Networks (CNNs) when data is scarce. Given the constraints of a small training set for the downstream task, we posit that incorporating a stronger inductive bias could improve the adaptation performance. In light of this, we propose Circular Convolution Adaptation (C3A), featuring the circular convolution operator Bamieh (2018).
Figure 2: Overview of LoRA (A) and our C3A (B,C) method. In LoRA, only low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ are trained and the delta weight is represented by their product (i.e., $\Delta\mathbf{W}=\mathbf{BA}$). The total trainable parameter number is $r(d_{1}+d_{2})$, which is assosiated with the rank of the delta weight. In C3A, circular convolution kernels $\Delta\mathbf{w}$ are tuned to adapt to the downstream task and the delta weight is represented by the (block-)circular matrix they construct (i.e., $\Delta\mathbf{W}=\mathcal{C}_{\mathrm{(blk)}}(\Delta\mathbf{w})$). The total trainable parameter count is $\frac{d_{1}d_{2}}{b}$, which disentangles with the rank of the delta weight. Here, $b$ is the block size of the block-circular matrix and it should be a common divisor (CD) of $d_{1}$ and $d_{2}$.
Circular convolution has been famous in signal processing Li et al. (2020) and cryptography Dworkin et al. (2001) for its distinguished efficiency and compactness. This operator can be equivalently represented as multiplication by a circulant matrix, which offers rank flexibility that is unconstrained by the number of trainable parameters. Moreover, leveraging the Fast Fourier Transform (FFT), C3A ensures better time and memory efficiency than directly multiplying the circulant matrix Bamieh (2018), rendering it competitive with LoRA from the efficiency aspect. In a word, circular convolution is a promising option for escaping the rank curse of LoRA at minimal cost. We summarize our contribution as follows:
@@
① We introduce C3A, a novel approach for PEFT using convolution. This method leverages the circular convolution operation and its equivalent circulant matrix to provide a flexible rank, which is free of linear constraint by the number of trainable parameters, for the delta matrix.
@@
④ We validate C3A through comprehensive fine-tuning experiments across diverse tasks including natural language understanding, instruction tuning and image classification. Experiments demonstrate C3A’s outstanding accuracy and memory merits compared to existing state-of-the-art methods.
@@ 2.1 Parameter-Efficient Fine-Tuning
Research on Parameter-Efficient Fine-Tuning (PEFT) has generally progressed along three main directions. The first direction involves partially updating the pre-trained neural network (e.g., the body Oh et al. (2020) or the biases Zaken et al. (2021)). Traditional methods relied on hand-crafted heuristics Han et al. (2015); Oh et al. (2020) to identify which parameters are crucial and should be fine-tuned. More advanced approaches employ optimization techniques Guo et al. (2020); Xu et al. (2021); Fu et al. (2023). For example, Guo et al. (2020) reformulated such a discrete optimization problem into a continuous one by employing Bernoulli masks and the Gumbel-softmax approximation.
The second direction emerged to maintain the integrity of the pre-trained model while enabling a high degree of parameter sharing through adapter-based methods He et al. (2021); Rebuffi et al. (2017); Rücklé et al. (2020); Liu et al. (2022). These works focus on integrating additional modules, termed adapters, to fit the downstream task, effectively decoupling the pre-trained model parameters from those specific to the downstream task. Prompt Tuning Brown et al. (2020); Gao et al. (2020); Chen et al. (2023); Zhang et al. (2024a) and Prefix Tuning Li & Liang (2021) also fall into this category, despite ignoring potential semantic meanings.
The final direction is characterized by delta-weight-based methods, such as Low-Rank Adaptation (LoRA) Hu et al. (2021) and Orthogonal Fine-tuning (OFT) Qiu et al. (2023). These methods bridge the gap between the pre-trained model and the downstream task by adaptive delta weights, which are stored separately while used in combination with the pre-trained weights. This unique design enables disentanglement of the pretrained and downstream-specific weights. Namely, it achieves parameter sharing and preserves the ability to integrate without additional inference cost. LoRA models the delta-weights by an additive matrix while OFT does it by a multiplicative one. To further improve either parameter efficiency or performance, many variants has been proposed for both of the methods Kopiczko et al. (2023); Liu et al. (2024; 2023); Yuan et al. (2024). However, these methods can hardly achieve high parameter efficiency and performance without incurring heavy computation and memory usage.
@@ 2.2 Circular Convolution
Despite its success in small neural networks such as LeNet Cheng et al. (2015), circular convolution has not demonstrated lossless performance in modern large foundational models (LFMs) or even in their base architecture, the transformer. This limitation may be attributed to the conflict between its high intrinsic bias (i.e., the circulant pattern) and the vast amount of data required for training LFMs. Conversely, when fine-tuning LFMs, it is often impractical to collect as much data as needed for training from scratch. In such scenarios, the intrinsic bias of circular convolution could potentially serve as a regularization mechanism, thereby benefiting the optimization process of fine-tuning.
@@ 3.1 Notations
The adapted weight matrix, the original weight matrix, and the delta matrix are denoted by $\mathbf{W}$, $\mathbf{W}_{0}$, and $\Delta\mathbf{W}$, respectively ($\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$). The activation vector of the previous layer is denoted by $\mathbf{x}\in\mathbb{R}^{d_{2}}$. The post-transformation vector is $\mathbf{z}$, where $\mathbf{z}=\mathbf{W}\mathbf{x}\in\mathbb{R}^{d_{1}}$, and the incremental part is denoted by $\Delta\mathbf{z}$, where $\Delta\mathbf{z}=\Delta\mathbf{W}\mathbf{x}\in\mathbb{R}^{d_{1}}$. The matrices $\mathbf{A}$ and $\mathbf{B}$ are low-rank matrices introduced by LoRA to represent $\Delta\mathbf{W}$, with $r$ being their rank. $r_{v}$ specifies the rank of the random projection matrix used in VeRA. The circular convolution kernel of C3A is denoted by $\Delta\mathbf{w}$ and the circular convolution operator by $\star$. The loss function is represented by $\mathcal{L}$. The Fast Fourier Transform and its inverse are denoted by $\mathrm{FFT}$ and $\mathrm{iFFT}$, respectively. The Hadamard product is denoted by $\circ$.
@@ 3.2 Circular Convolution
Firstly, for simplicity, we assume $d_{1}=d_{2}=d$ and $\Delta\mathbf{w}\in\mathbb{R}^{d}$. The circular convolution operator is defined as $\Delta\mathbf{z}=\Delta\mathbf{w}\star\mathbf{x}=\mathcal{C}(\Delta\mathbf{w})\mathbf{x}$, where $\mathcal{C}(\cdot)$ is a function which takes a vector and outputs the corresponding circulant matrix. Concretely, the first row of $\mathcal{C}(\Delta\mathbf{w})$ is $\Delta\mathbf{w}$ and the following rows are equal to the row above them periodically shifted to the right by one element. In math,
$\mathcal{C}(\Delta\mathbf{w})=\begin{bmatrix}\Delta w_{1}&\Delta w_{2}&\cdots&\Delta w_{d-1}&\Delta w_{d}\\ \Delta w_{d}&\Delta w_{1}&\cdots&\Delta w_{d-2}&\Delta w_{d-1}\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ \Delta w_{3}&\Delta w_{4}&\cdots&\Delta w_{1}&\Delta w_{2}\\ \Delta w_{2}&\Delta w_{3}&\cdots&\Delta w_{d}&\Delta w_{1}\\ \end{bmatrix}.$
Theoretically, the rank of $\mathcal{C}(\Delta\mathbf{w})$ is given by $d-\mathrm{Deg}(\mathrm{gcd}(f(x),x^{d}-1))$ Ingleton (1956), where $\mathrm{Deg}(\cdot)$ denotes the degree of a polynomial, $f(x)$ is the polynomial associated with $\Delta\mathbf{w}$ (i.e., $f(x)=\sum_{i=1}^{d}\Delta w_{i}x^{i-1}$), and $\mathrm{gcd}(\cdot)$ represents the greatest common divisor. Consequently, the theoretical upper bound on the rank of $\mathcal{C}(\Delta\mathbf{w})$ is $d$. By learning $\Delta\mathbf{w}$ in the $\mathbb{R}^{n}$ oracle, C3A automatically achieves dynamic rank selection, which is not linearly constrained by the number of learnable parameters, unlike LoRA.
To achieve high efficiency, we leverage the beautiful circulant structure of $\mathcal{C}(\Delta\mathbf{w})$, which makes it diagonalizable by the Fourier basis ($\mathbf{F}$) . In math, it can be described as $\mathcal{C}(\Delta\mathbf{w})=\mathbf{F}\frac{\Lambda}{d}\mathbf{F}^{-1}$ Golub & Van Loan (1996), where $\Lambda$ is its eigenvalues and can be calculated by a Fourier transform of the first row (i.e., $\Lambda=\mathrm{diag}(\mathbf{F}\Delta\mathbf{w})$). Therefore, we can calculate $\Delta\mathbf{w}\star\mathbf{x}$ as
@@ 3.3 Backpropagation
Since $\Delta\mathbf{z}=\mathcal{C}(\Delta\mathbf{w})\mathbf{x}$, it follows that $\frac{\partial\Delta\mathbf{z}}{\partial\mathbf{x}}=\mathcal{C}(\Delta\mathbf{w})$. For $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}$, we note that the circular convolution operation is commutative (i.e., $\mathcal{C}(\Delta\mathbf{w})\mathbf{x}=\mathcal{C}(\mathbf{x})\Delta\mathbf{w}$), thus $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}=\mathcal{C}(\mathbf{x})$. Substituting these results into Equation 2, we obtain:
@@ 3.4 Block-Circular Convolution
Despite the elegance and efficiency of the circular convolution operator, it has 2 fundamental limitations due to the requirement that the convolution kernel must be the same size as the activation vector: ❶ It cannot be applied to non-square weight matrices.The number of learnable parameters cannot be altered. The first limitation restricts its application in scenarios such as fine-tuning a LLaMA3-8B model, where the weight matrix sizes include $4096\times 1024$. The second limitation reduces the flexibility of C3A, making it challenging to handle complex downstream tasks that require a larger number of learnable parameters. To address these limitations, we resort to block-circular convolution. By splitting the activation vector $\mathbf{x}$ and the post-transformation vector $\mathbf{z}$ into blocks of the same size, we can assign unique convolution kernels to each pair of these blocks. To be concrete,
$\mathbf{x}=\begin{bmatrix}\mathbf{x}_{1}&\mathbf{x}_{2}&\cdots&\mathbf{x}_{\frac{d_{2}}{b}}\end{bmatrix},\qquad\Delta\mathbf{z}=\begin{bmatrix}\Delta\mathbf{z}_{1}&\Delta\mathbf{z}_{2}&\cdots&\Delta\mathbf{z}_{\frac{d_{1}}{b}}\end{bmatrix},$
where $b$ is the block size and $b$ need to be a common divisor of $d_{1}$ and $d_{2}$. We will need $\frac{d_{1}d_{2}}{b^{2}}$ convolution kernels to densely connect these blocks, which can be expressed in math as
$\Delta\mathbf{z}_{i}=\sum_{j=1}^{\frac{d_{2}}{b}}\Delta\mathbf{w}_{ij}\star\mathbf{x}_{j},i\in\{1,2,\cdots,\frac{d_{1}}{b}\}.$
This calculation can be represented by a block-circular matrix:
$\Delta\mathbf{z}=\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\mathbf{x},\qquad\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})=\begin{bmatrix}\mathcal{C}(\Delta\mathbf{w}_{11})&\mathcal{C}(\Delta\mathbf{w}_{12})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{1\frac{d_{2}}{b}})\\ \mathcal{C}(\Delta\mathbf{w}_{21})&\mathcal{C}(\Delta\mathbf{w}_{22})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{2\frac{d_{2}}{b}})\\ \cdots&\cdots&\cdots&\cdots\\ \mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}1})&\mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}2})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}\frac{d_{2}}{b}})\end{bmatrix}.$
In this case, $\Delta\mathbf{w}_{ij}\in\mathbb{R}^{b}$ and we have $\frac{d_{1}d_{2}}{b^{2}}b=\frac{d_{1}d_{2}}{b}$ learnable parameters. In fact, $b$ acts as a hyperparameter controlling the number of learnable parameters, just like $r$ for LoRA. However, we highlight that while $r$ controls the rank of the delta matrix and learnable parameter number at the same time, $b$ only affects the latter. Disentangling these attributes enables higher adaptation flexibility and potential better results.